OpenStudy (anonymous):

[TRIG] What part of a complete revolution is an angle of 45 degrees?

jimthompson5910 (jim_thompson5910):

what is a complete revolution

OpenStudy (anonymous):

360 degrees

jimthompson5910 (jim_thompson5910):

so to answer the question, you need to reduce 45/360

jimthompson5910 (jim_thompson5910):

for instance, if the angle was 180 (instead of 45), then it would be 180/360 = 1/2 of a full revolution

OpenStudy (anonymous):

and i get 8, oh i see

jimthompson5910 (jim_thompson5910):

1/8 actually

jimthompson5910 (jim_thompson5910):

45/360 = 1/8

jimthompson5910 (jim_thompson5910):

so 45 degrees is 1/8 of a full rotation

OpenStudy (anonymous):

45/360 = 0.125 though..

jimthompson5910 (jim_thompson5910):

1/8 = 0.125

jimthompson5910 (jim_thompson5910):

that's why

OpenStudy (anonymous):

for future reference how would i translate the decimal to fraction form?

jimthompson5910 (jim_thompson5910):

0.125 = 125/1000 since the 5 is in the thousandths place

jimthompson5910 (jim_thompson5910):

125/1000 would then reduce to 1/8 after dividing both parts by 125 (the GCF)

jimthompson5910 (jim_thompson5910):

this trick only works with decimals that terminate

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

it won't work with repeating decimals

jimthompson5910 (jim_thompson5910):

yw